# Rethinking Oversmoothing in Graph Neural Networks: A Rank-Based Perspective

### v2

@@ A.3 An illustrative example
Appendix B Additional Empirical Results on Real Dataset
LReLU + GCN $r^{*}_{\mathrm{ER}}=1.450100,\quad r^{*}_{\mathrm{NR}}=1.207918$ Tanh + GCN $r^{*}_{\mathrm{ER}}=1.690520,\quad r^{*}_{\mathrm{NR}}=1.007135$
@@ Appendix B Additional Empirical Results on Real Dataset
LReLU + GCN + Bias $r^{*}_{\mathrm{ER}}=1.454981,\quad r^{*}_{\mathrm{NR}}=1.020319$ Tanh + GCN + Bias $r^{*}_{\mathrm{ER}}=1.858602,\quad r^{*}_{\mathrm{NR}}=1.021634$
@@
LReLU + GCN + LayerNorm $r^{*}_{\mathrm{ER}}=1.017424,\quad r^{*}_{\mathrm{NR}}=1.000101$ Tanh + GCN + LayerNorm $r^{*}_{\mathrm{ER}}=1.528688,\quad r^{*}_{\mathrm{NR}}=1.003682$
@@
LReLU + GAT $r^{*}_{\mathrm{ER}}=1.295623,\quad r^{*}_{\mathrm{NR}}=1.079250$ Tanh + GAT $r^{*}_{\mathrm{ER}}=1.002191,\quad r^{*}_{\mathrm{NR}}=1.000006$
@@
LReLU + GCN $r^{*}_{\mathrm{ER}}=1.750563,\quad r^{*}_{\mathrm{NR}}=1.142120$ Tanh + GCN $r^{*}_{\mathrm{ER}}=1.436012,\quad r^{*}_{\mathrm{NR}}=1.003035$
@@
LReLU + GCN + Bias $r^{*}_{\mathrm{ER}}=1.526297,\quad r^{*}_{\mathrm{NR}}=1.045451$ Tanh + GCN + Bias $r^{*}_{\mathrm{ER}}=1.752850,\quad r^{*}_{\mathrm{NR}}=1.007417$
@@
LReLU + GCN + LayerNorm $r^{*}_{\mathrm{ER}}=1.006405,\quad r^{*}_{\mathrm{NR}}=1.000009$ Tanh + GCN + LayerNorm $r^{*}_{\mathrm{ER}}=1.473417,\quad r^{*}_{\mathrm{NR}}=1.002555$
@@
LReLU + GAT $r^{*}_{\mathrm{ER}}=1.541711,\quad r^{*}_{\mathrm{NR}}=1.148837$ Tanh + GAT $r^{*}_{\mathrm{ER}}=1.003797,\quad r^{*}_{\mathrm{NR}}=1.000005$
@@
LReLU + GCN $r^{*}_{\mathrm{ER}}=1.420913,\quad r^{*}_{\mathrm{NR}}=1.085279$ Tanh + GCN $r^{*}_{\mathrm{ER}}=1.838982,\quad r^{*}_{\mathrm{NR}}=1.014851$
@@
LReLU + GCN + Bias $r^{*}_{\mathrm{ER}}=1.450100,\quad r^{*}_{\mathrm{NR}}=1.207918$ Tanh + GCN + Bias $r^{*}_{\mathrm{ER}}=1.443576,\quad r^{*}_{\mathrm{NR}}=1.018361$
@@
LReLU + GCN + LayerNorm $r^{*}_{\mathrm{ER}}=1.002088,\quad r^{*}_{\mathrm{NR}}=1.000074$ Tanh + GCN + LayerNorm $r^{*}_{\mathrm{ER}}=1.668289,\quad r^{*}_{\mathrm{NR}}=1.009275$
@@
LReLU + GAT $r^{*}_{\mathrm{ER}}=1.577163\quad r^{*}_{\mathrm{NR}}=1.119213$ Tanh + GAT $r^{*}_{\mathrm{ER}}=1.016096,\quad r^{*}_{\mathrm{NR}}=1.000019$

### v3

@@ A.3 An illustrative example
Appendix B Limitations
Although a small effective rank or numerical rank indicates oversmoothing and can be subsequently linked to the underperformance of GNNs, a large effective rank or numerical rank does not necessarily correspond to a good network performance. Prior study has suggested some degree of smoothing can be beneficial [16], and as an extreme example, features sampled from a uniform distribution and with randomly assigned labels would almost surely have a large effective rank, but they cannot be classified accurately due to the lack of any underlying pattern. We note that this limitation is not specific to the two relaxed rank measures, as the same argument is directly applicable to all other oversmoothing metrics.
Consequently, as shown in Table 4, when additional components are used to (partially) alleviate oversmoothing, particularly when residual connections are used, the accuracy ratio may remain large over the layers, and all oversmoothing metrics correlate poorly with the accuracy of GNNs. This in-turn suggests these oversmoothing metrics become less informative as the oversmoothing problem is mitigated or alleviated.
Appendix C Computational Complexity Analysis
Let $N$, $D$, $E$ the number of nodes, features and edges of a graph $\mathcal{G}$, the computational cost of the Dirichlet energy is $O(E\times D)$, while the cost of the Projection Energy is $O(N\times D)$. In contrast, most of the computational cost for the numerical rank and the effective rank is given by the computation of the spectral radius of $X$, and of all the singular values of $X$ for the effective rank, respectively.
Standard results show that it is always possible to compute the full singular value decomposition (SVD), and subsequently the spectral radius, in $O(N\times D\times\min\{N,D\})$. However, in typical cases, the latter cost can be drastically reduced using different strategies. Firstly, in the case of numerical rank, the computational cost of the 2-norm is generally much smaller than the cost of the full SVD. Indeed, using Lanczos or power methods to compute it, the cost scales as $O(N\times D)$, and the methods converge typically very fast. In addition, both the effective rank and the numerical rank can be efficiently controlled by computing only the $k$-largest singular values of $X$. In particular, a truncated SVD containing the $k$-largest singular values can be computed in $O(N\times D\times k)$ using either deterministic algorithms or randomized SVD methods.
In general, the computational cost of metrics to quantify oversmoothing is marginal compared to the cost of training. In production, measuring the emergence of oversmoothing is done by training the model and checking the performance on the validation and test sets as the number of layers changes. The cost of additionally computing effective or numerical ranks is marginal. Moreover, we note that computing the metrics on a subsection of a graph can be sufficiently informative, and as a consequence, most oversmoothing metrics studied in this paper can be computed in less than $10~{}ms$ for a graph (or a subsection of it) with less than 2000 nodes.
Appendix D Additional Experimental Results with Synthetic Weights
In this section, we conduct an asymptotic ablation study using randomly sampled (synthetic) untrained weights. The aim of these experiments is twofold:
  • to demonstrate that similar untrained asymptotic experiments are inherently unrealistic, despite being extensively used in the literature [44, 35, 34, 45, 29, 42], as they fail to reliably capture oversmoothing in shallower GNNs where the performance degradation occurs.
  • to examine the convergence properties of different oversmoothing metrics with weight size control and to empirically validate Theorems 5.1 and 5.4.
We construct a 10-node Barabasi-Albert graph with each node having 32 features. The weights are either an identity matrix or randomly sampled at each layer from a uniform distribution $\mathcal{U}(0,s)$, where $s$ depends on the settings: small weights ($s=0.05$) lead to an exponentially decaying $\|X^{(l)}\|_{F}$, and large weights ($s=0.1$) lead to an exploding $\|X^{(l)}\|_{F}$ for uncapped activation functions. For identity weights, $\|X^{(l)}\|_{F}$ is roughly constant (LReLU) or slowly decaying (Tanh). The feature initialization $X^{(0)}$ is sampled from $\mathcal{U}(0,1)$, and is iterated over 300 layers.
In this asymptotic synthetic setting, as presented in Table 2, the normalized $E_{\mathrm{Dir}}$ and $E_{\mathrm{Proj}}$ exhibit decay patterns similar to those of the effective rank and numerical rank, suggesting these metrics are equally sensitive to asymptotic rank collapse. However, this behaviour stands in stark contrast to results on trained networks presented in Tables 1 and 3 and Appendix G, where the normalized $E_{\mathrm{Dir}}$ and $E_{\mathrm{Proj}}$ often fail to detect oversmoothing.
Moreover, these asymptotic results validate Theorems 5.1 and 5.4, showing that the numerical rank converges to one for GCN + LReLU and GAT + any subhomogeneous activation functions. Without making any additional assumption on the normalization of the adjacency matrix, the effective rank and numerical rank do not generally decay to one when subhomogeneous activation functions, e.g. Tanh, are used in GCNs.
Architecture $E_{\mathrm{Dir}}$ $E_{\mathrm{Proj}}$ MAD Erank NumRank
standard normalized standard normalized
GCN+LReLU+identity weights
GCN+Tanh+identity weights
GAT+LReLU+identity weights
GAT+Tanh+identity weights
GCN+LReLU+small weights
GCN+Tanh+small weights
GAT+LReLU+small weights
GAT+Tanh+small weights
GCN+LReLU+large weights
GCN+Tanh+large weights
GAT+LReLU+large weights
GAT+Tanh+large weights
Table 2: Additional results on very deep (300 layers) synthetic networks with randomly sampled weights. For Erank and NumRank, we subtract 1 so that both metrics converge to zero. ✓ indicates a decay of the corresponding metric to zero, ✗ indicates otherwise. Note that GAT has similar asymptotic behaviour to GCN with adjacency normalization $D^{-1}\widetilde{A}$.
Appendix E Additional Experimental Results on Activation Functions and Different Datasets
We extend Table 1 to subhomogenous Tanh activation function and a few additional datasets. The experimental setting is consistent with that of Section 6.
Dataset Architecture $E_{\mathrm{Dir}}$ $E_{\mathrm{Proj}}$ MAD Erank NumRank Accuracy ratio
Standard Normalized Standard Normalized
Cora GCN+LReLU -0.7871 0.6644 -0.8106 -0.8309 -0.2460 0.9724 0.5885 0.2693
GCN+Tanh 0.5243 0.9403 0.8610 0.9768 0.9734 0.9923 0.9784 0.1937
GAT+LReLU -0.9189 0.6703 -0.9469 -0.6054 0.8251 0.9722 0.7612 0.2493
GAT+Tanh 0.8300 0.8501 0.8676 0.9066 0.8603 0.9600 0.9515 0.1900
Citeseer GCN+LReLU -0.8442 0.4350 -0.8913 -0.8667 -0.7169 0.9700 0.6795 0.4380
GCN+Tanh 0.3420 0.8957 0.4631 0.9045 0.9605 0.9906 0.9457 0.3509
GAT+LReLU -0.9576 0.0664 -0.9585 -0.9080 0.3722 0.9915 0.8047 0.4672
GAT+Tanh 0.9234 0.8949 0.8276 0.8997 0.9176 0.9287 0.9024 0.3045
Pubmed GCN+LReLU -0.9068 0.7006 -0.8508 -0.1109 0.6205 0.9464 0.9268 0.5225
GCN+Tanh -0.2330 0.2657 0.2029 0.9137 0.8745 0.9328 0.9940 0.3883
GAT+LReLU -0.8735 -0.3684 -0.8541 -0.4102 -0.3932 0.9270 0.9721 0.5564
GAT+Tanh 0.2977 0.8411 0.7160 0.9331 0.8546 0.9303 0.8551 0.4464
Squirrel GCN+LReLU -0.7774 0.4171 -0.7602 -0.3258 -0.8247 0.6316 0.9582 0.8457
GCN+Tanh 0.6026 0.7736 0.1689 0.9377 0.7727 0.9680 0.9837 0.8152
GAT+LReLU -0.6864 -0.5503 -0.7364 -0.7253 0.5002 0.8538 0.6840 0.7533
GAT+Tanh -0.3606 -0.6557 -0.0363 0.8714 -0.7033 0.8911 0.8861 0.9103
Chameleon GCN+LReLU -0.9223 0.1504 -0.9163 -0.8201 -0.8809 0.9387 0.9014 0.6195
GCN+Tanh 0.2742 0.8796 -0.2541 0.8492 0.9272 0.9869 0.9841 0.7093
GAT+LReLU -0.8721 0.1942 -0.9089 -0.8234 0.2803 0.9446 0.8799 0.6332
GAT+Tanh 0.4090 0.5699 0.0743 0.8230 0.6006 0.9613 0.9143 0.7052
Amazon Ratings GCN+LReLU -0.9297 0.8809 -0.9079 -0.3423 0.9201 0.9301 0.8049 0.8562
GCN+Tanh -0.6960 -0.6289 -0.6910 0.8576 0.9423 0.9871 0.9327 0.8574
GAT+LReLU -0.9388 0.5277 -0.9089 -0.1617 0.6545 0.9248 0.8764 0.8384
GAT+Tanh 0.8354 0.8507 -0.6595 0.9245 0.9418 0.8954 0.8883 0.8382
Roman Empire GCN+LReLU -0.5635 0.7703 -0.6772 0.2575 0.6420 0.5833 0.5368 0.3891
GCN+Tanh 0.8018 0.9225 -0.6808 0.8359 0.8124 0.8570 0.8090 0.4067
GAT+LReLU -0.9174 0.5390 -0.9407 0.1868 0.7582 0.7221 0.8722 0.3705
GAT+Tanh 0.5767 0.5844 -0.4716 0.8819 0.7263 0.7332 0.8589 0.3652
OGB-Arxiv GCN+LReLU 0.7738 0.9194 0.5740 -0.2738 0.2822 0.9682 0.9091 0.0957
GCN+Tanh -0.6409 0.7041 -0.6808 0.9494 0.9487 0.9900 0.9876 0.1204
GAT+LReLU -0.4097 0.9439 -0.7230 0.8985 0.8492 0.7740 0.9781 0.2310
GAT+Tanh 0.7834 0.7896 -0.9277 0.9393 0.8222 0.7671 0.7861 0.2376
Average correlation -0.1956 0.5137 -0.3886 0.2669 0.4960 0.9007 0.8684
Table 3: Additional correlation coefficient results on homophilic (Cora, Citeseer, Pubmed), heterophilic (Squirrel, Chameleon, Amazon Ratings, Roman Empire) and large-scale (OGB-Arxiv) dataset.
Appendix F Additional Experimental Results with Different Network Components
Prior literature indicates that when adding additional components, such as bias [33] or residual terms [36], the Dirichlet energy does not decay. Therefore, we compare the correlation coefficients in Table 4 between existing oversmoothing metrics when additional components are added, such as bias, LayerNorm, BatchNorm, PairNorm [46], DropEdge [28, 15] and residual connections [36].
All experiments follow the setup described in Section 6. In addition, DropEdge has a probability of 0.5 in removing each edge at each layer. The residual connection is implemented as follows
$X^{(l+1)}=\sigma(AX^{(l)}W_{1}^{(l)})+X^{(0)}W_{2}^{(l)}$
Table 4 demonstrates that both the effective rank and numerical rank achieve a higher average correlation with the classification accuracy than alternative metrics. This finding confirms their superior consistency in detecting oversmoothing across a variety of architectural variants.
Furthermore, we note that when oversmoothing is effectively alleviated, e.g. when residual connection is used, all metrics have a poor correlation with the classification accuracy. This generic limitation is discussed in Appendix B.
Standard Normalized Standard Normalized
GCN+LReLU+Bias -0.9300 0.7505 -0.9414 -0.3642 -0.2828 0.9847 0.7833 0.2110
GCN+Tanh+Bias 0.4086 0.8950 0.7479 0.9026 0.8996 0.9926 0.9814 0.2133
GCN+LReLU+LayerNorm -0.9132 0.8445 -0.9424 0.5591 0.9753 0.9736 0.9736 0.4882
GCN+Tanh+LayerNorm -0.7119 0.2560 -0.2069 0.9716 0.8695 0.9576 0.9684 0.1886
GCN+LReLU+BatchNorm 0.8789 0.7300 0.8872 0.7175 0.5094 0.6005 0.6577 0.8761
GCN+Tanh+BatchNorm -0.6033 0.2883 -0.6587 0.7717 0.4984 0.7377 0.7038 0.8157
GCN+LReLU+PairNorm -0.8165 0.3731 -0.8106 0.3885 0.4850 0.5597 0.6244 0.8556
GCN+Tanh+PairNorm -0.5963 0.0346 -0.5401 -0.3358 -0.0631 0.9838 0.9298 0.3123
GCN+LReLU+DropEdge -0.7497 0.7185 -0.7939 -0.7974 -0.4619 0.9720 0.5782 0.2515
GCN+Tanh+DropEdge 0.2319 0.8388 0.5872 0.8887 0.8763 0.9934 0.9558 0.2096
GCN+LReLU+Residual -0.0857 -0.3072 -0.0559 -0.1611 -0.1656 -0.2466 -0.1296 1.0046
GCN+Tanh+Residual -0.6751 -0.7406 -0.3031 -0.4019 -0.6043 -0.5425 -0.4897 1.0122
Average correlation -0.3801 0.3901 -0.2525 0.2616 0.2946 0.6638 0.6280
Table 4: Additional correlation coefficient results on GCNs with different network components.
Appendix G Metric Behaviour Examples on Cora, Citeseer and Pubmed
We extend Figure 2 to additional datasets and network components. The experimental setting is consistent with that of Section 6.
GCN + LReLU $r^{*}_{\mathrm{ER}}=1.450100,\quad r^{*}_{\mathrm{NR}}=1.207918$ GCN + Tanh $r^{*}_{\mathrm{ER}}=1.690520,\quad r^{*}_{\mathrm{NR}}=1.007135$
@@ Appendix G Metric Behaviour Examples on Cora, Citeseer and Pubmed
GCN + LReLU + Bias $r^{*}_{\mathrm{ER}}=1.454981,\quad r^{*}_{\mathrm{NR}}=1.020319$ GCN + Tanh + Bias $r^{*}_{\mathrm{ER}}=1.858602,\quad r^{*}_{\mathrm{NR}}=1.021634$
@@
GCN + LReLU + LayerNorm $r^{*}_{\mathrm{ER}}=1.017424,\quad r^{*}_{\mathrm{NR}}=1.000101$ GCN + Tanh + LayerNorm $r^{*}_{\mathrm{ER}}=1.528688,\quad r^{*}_{\mathrm{NR}}=1.003682$
@@
GAT + LReLU $r^{*}_{\mathrm{ER}}=1.295623,\quad r^{*}_{\mathrm{NR}}=1.079250$ GAT + Tanh $r^{*}_{\mathrm{ER}}=1.002191,\quad r^{*}_{\mathrm{NR}}=1.000006$
@@
GCN + LReLU $r^{*}_{\mathrm{ER}}=1.750563,\quad r^{*}_{\mathrm{NR}}=1.142120$ GCN + Tanh $r^{*}_{\mathrm{ER}}=1.436012,\quad r^{*}_{\mathrm{NR}}=1.003035$
@@
GCN + LReLU + Bias $r^{*}_{\mathrm{ER}}=1.526297,\quad r^{*}_{\mathrm{NR}}=1.045451$ GCN + Tanh + Bias $r^{*}_{\mathrm{ER}}=1.752850,\quad r^{*}_{\mathrm{NR}}=1.007417$
@@
GCN + LReLU + LayerNorm $r^{*}_{\mathrm{ER}}=1.006405,\quad r^{*}_{\mathrm{NR}}=1.000009$ GCN + Tanh + LayerNorm $r^{*}_{\mathrm{ER}}=1.473417,\quad r^{*}_{\mathrm{NR}}=1.002555$
@@
GAT + LReLU $r^{*}_{\mathrm{ER}}=1.541711,\quad r^{*}_{\mathrm{NR}}=1.148837$ GAT + Tanh $r^{*}_{\mathrm{ER}}=1.003797,\quad r^{*}_{\mathrm{NR}}=1.000005$
@@
GCN + LReLU $r^{*}_{\mathrm{ER}}=1.420913,\quad r^{*}_{\mathrm{NR}}=1.085279$ GCN + Tanh $r^{*}_{\mathrm{ER}}=1.838982,\quad r^{*}_{\mathrm{NR}}=1.014851$
@@
GCN + LReLU + Bias $r^{*}_{\mathrm{ER}}=1.450100,\quad r^{*}_{\mathrm{NR}}=1.207918$ GCN + Tanh + Bias $r^{*}_{\mathrm{ER}}=1.443576,\quad r^{*}_{\mathrm{NR}}=1.018361$
@@
GCN + LReLU + LayerNorm $r^{*}_{\mathrm{ER}}=1.002088,\quad r^{*}_{\mathrm{NR}}=1.000074$ GCN + Tanh + LayerNorm $r^{*}_{\mathrm{ER}}=1.668289,\quad r^{*}_{\mathrm{NR}}=1.009275$
@@
GAT + LReLU $r^{*}_{\mathrm{ER}}=1.577163\quad r^{*}_{\mathrm{NR}}=1.119213$ GAT + Tanh $r^{*}_{\mathrm{ER}}=1.016096,\quad r^{*}_{\mathrm{NR}}=1.000019$